Operation Summary for January 15, 2010 - January 18, 2010 |
The Integrated Luminosity for the period from 1/11/10 to 1/18/10 was 48.57 pb-1. Friday Jan. 15 The day shift began with Operations monitoring the Tevatron (TeV) store 7507 (110.42E30), the stash (328.4E10), and the Antiproton Source stack (7.5E10), with Linac personnel working on a RF station (LRF2), with Booster in access, with Main Injector (MI) operating normally, but with no beam, with the NuMI and MiniBooNE experiments off due to Proton Source work, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Linac resumed sending beam at 9:03 AM. Booster personnel completed their access and Operations had beam restored to Pbar and NuMI by 11:45 AM and with beam to MiniBooNE a short time later. Operations established beam to MTest (T977 MINERvA) at 12:50 PM. Store 7507 aborted at 2:25 PM due to a quadrupole power supply (BQ0) faulting off. EE Support personnel worked on the supply to the end of shift. During the evening shift, EE Support personnel completed their repairs on BQ0 at 4:35 PM. Operations put the TeV into HEP shot setup at 6:40 PM and established store 7509 at 8:05 PM with an average initial luminosity (AIL) of 252.43E30. Operations disabled the MTest beam at 6 PM. Saturday Jan. 16 The midnight shift began with Operations monitoring store 7509 (136.47E30) and the stash (172.0E10), with Pbar stacking (21.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operators search and secured the MT6 enclosure. Operations established beam to MTest at 4:20 AM. During the day shift, a MI injection kicker power supply (KPS1B) tripped off at 8 AM; EE Support personnel came in and replaced the resonant charger power supply. MI resumed sending beam by 1:10 PM. During the evening shift, Operations terminated store 7509 at 4:45 PM and put the TeV into shot setup. Operations established store 7510 at 6:30 PM with an AIL of 271.3E30. Sunday Jan. 17 The midnight shift began with Operations monitoring store 7510 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operators searched and secured MT6. Operations established beam to MTest at 4:20 AM. During the day shift, Operations terminated store 7510 at 10:15 AM and then put the TeV into shot setup. Operations established store 7511 at 11:55 AM with an AIL of 270.2E30. During the evening shift, Operations disabled the MTest beam at 6 PM. Monday Jan. 18 The midnight shift began with Operations monitoring store 7511 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations terminated store 7511 at 4:25 AM and then put the TeV into shot setup. Operations established store 7512 at 5:55 AM with an AIL of 245.91E30. Operations established beam to MTest at 5:54 AM. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 02/03/2010 email Fermilab |
| Security, Privacy, Legal | |